ERM operational status:
The station operator visited the station and replaced both Q330s with new Q330s with upgraded firmware.
The station is back online as of day 2021,074 after the station operator visited the station and reconnected the LAN cable to the Netgear switch in the recording room.
The station is down as of 2021,063 due to loss of communication between the Q330 digitizers and station computers. IDA engineers are working with the station operator to resolve this issue.
The station is back online as of day 2020,189 (July7) after the station operator visited the station and replaced the station computers. Another station visit will be planned to resolve other power issues at the station.
The station is down due to power and station computer issues. When cleared to visit the station, the station operator will investigate the power issues, replace batteries, breaker, and station computers.
The station operator visited the station and reestablished the vacuum and improved the data quality for the M2166 seismometer. There was increased noise from the M2166 seismometer since an offset on day 2015,244 due to loss of vacuum.
An IDA engineer visited the station to upgrade the power system, replace the STS-1 sensors with new Metrozet M2166 sensors, switch the Setra microbarometer connection to the ERM-10 digitizer, and replace the station computers.
IDA station ERM has been upgraded to a Quanterra Q330HR system. Sensors include STS-1 and STS-2.5 seismometers, an episensor strong motion sensor, and a Setra barometer. The north-south STS-1 was replaced during the upgrade. However, the north-south and vertical STS-1 channels are still showing long period noise. The base plates for the sensors are the suspected cause for this noise.
IDA engineers are enroute to upgrade the station to a Q330 system.
IDA station ERM is back online as of day 2012,233-01:13 after the station operator cycled power to the DAS.
IDA station ERM is down since day 2012,225-03:35 due to problems with the DAS. The station operator has been requested to cycle power to the DAS.
The station is back on-line as of day 2012,187-08:42 after the station operator visited the station and cycled power to the digitizer.
The station is down as of day 2012,177-22:37 due to problems with the DAS that are yet to be determined.
Telemetry to the station was restored after the station operator replaced the router.
Telemetry to IDA station ERM has been down since day 2012,127. IDA engineers are working with the station operator to restore the link to the station.
The North-South channels from the STS-1 seismometer are non-seismic as of day 2011,261-22:43.
The North-South channels from the STS-1 seismometer were non-seismic from day 2010,287-13:59 to 2010,306-12:39.
The station is back on-line as of day 2009,301-18:51, after the station operator visited the station and restored power. An IDA engineer then restored data acquisition after cycling power to the DAS and ARS.
The station is down as of day 2009,281-10:59 due to a power outage caused by a storm. The station operator notified he will visit the station to investigate.
The station is back online as of about day 2007,247-02:38 after the station operator reploced the ISP.
The station is down due to problems with the ISP that are yet to be determined.
The station operator replaced a faulty GPS clock with a new one shipped from UCSD. The clock was in an unlocked state since day 331-00:31. However, data remained continuous and data quality remains unaffected.
The vertical channels for the STS-1 seismometer are back as of about day 243-02:40 after an IDA engineer visited the site and replaced the STS-1 seismometer. The mass position for this seismometer is drifting and there appears to be a problem with the motor for this sensor. Also installed was an iboot to allow remote power cyles of the DAS.
The East-West channel from the STS-1 sensor is noisy since a DAS reset on 2006,116 (Apr. 25). A trip to the station by an IDA engineer is planned.
The station is back up after the station operator cycled power to the DAS.
The station is down due to communication problems between the DAS and the ISP.
The station is back up after the station operator rebooted the router.
The station is down as of Jan.12 for reasons yet to be determined.
The ARS was power cycled at around 2005:331-16:50 (Nov 27) and data acquisition resumed.
The station is down for reasons yet to be determined.
Station is back up after station operator reset the DAS and ARS and reconfigured connections to the STS-1 sensors.
Station is down as of day 346-10:54 (Dec 11). The ARS needs to be reset.
Station is operational. The vertical channels from the STS-1 sensor are still out.
Station operator replaced the broken GPS clock and NRTS computer. Telemetry is up after station operator initiated a high speed dedicated internet connection.
Station operator replaced DAT drive, reports problem with GPS receiver. The vertical channels from the STS-1 and FBA-23 sensors are non-seismic.
Telemetry down as of 12/29 due to problems connecting to modem.
NRTS data continuously acquired again as of day 226 (August 14). Intermittent NRTS coverage between days 208 and 226.
Telemetry down since 208-16:14(July 27,2003). Station operator requested to power cycle the data logger.
No NRTS data. ERM station operator power cycled data logger on 2003.05.14, ERM disk loop acquiring data again.
New EPROMs installed to correct DSP problem.
Recording at the station has been problematic recently. A telemetry check of the ARS on 2002.04.22 confirms recording is still normal. On 2002.04.13 the operator cycled power to data acquisition system. Operation returned to normal at that time.